OpenStudy (anonymous):

log2x=4 i need help.

OpenStudy (anonymous):

you want to find x?

OpenStudy (anonymous):

yes i believe so.

OpenStudy (anonymous):

is it \[\log_{2} x?\]

OpenStudy (anonymous):

or \[\log(2x)\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

if it is \[\log(2x)=4\] then exponentiate both sides i.e,\[e[\log(2x)]=e^4\] it becomes: \[2x=e^4\] \[x=\frac{ e^4 }{ 2 }\]